In a flight of 600 km, an aircraft was slowed down due to bad weather. Its average speed for the trip was reduced to 200km/hr and the time increased by 30 minutes. Find the duration of the flight.
Answer by Alan3354(69443)

In a flight of 600 km, an aircraft was slowed down due to bad weather. Its average speed for the trip was reduced to 200km/hr and the time increased by 30 minutes. Find the duration of the flight.
------------
600 = r*t = 200*(t + 0.5)
600 = 200t + 100
t = 2.5 hours